OpenStudy (anonymous):

Given the equation y - 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in standard form. -three fourthsx + y = 10 3x - 4y = -40 y = three fourthsx + 12 y = three fourthsx + 10

OpenStudy (anonymous):

All you have to do is solve for y. y-4 = 3/4(x+8) But first distribute the right side. y-4 = 3/4x + 6. y-4 = 3/4x + 10.